OpenStudy (anonymous):

What is the value of the expression 3^4 – 4^3?

OpenStudy (anonymous):

@oneinmillon

OpenStudy (campbell_st):

you need a calculator as its 3 x 3 x 3 x 3 - 4 x 4 x 4 =

OpenStudy (whpalmer4):

or just figure it out. \[3^4:\]\[3*3 = 9\]\[3*3*3 = 9*3=27\]\[3*3*3*3=9*3*3=27*3=81\] \[4^3:\]\[4*4=16\]\[4*4*4=16*4=64\] \[3^4-4^3=\]

OpenStudy (anonymous):

17?

OpenStudy (whpalmer4):

Yes!

OpenStudy (anonymous):

Which is 1024 written as a power of a number? A. 4^5 B. 256^4 C. 5^4 D. 512^2

OpenStudy (whpalmer4):

\[4^5= 4*4*4*4*4 = \]\[256^4=256*256*256*256 =\]\[5^4=5*5*5*5 =\]\[512^2=512*512=\] Should be easy enough to eliminate B and D from consideration, as they'll be way too big. As for C, if you multiply a counting number by 5, can you ever have a final digit other than 0 or 5?

OpenStudy (anonymous):

is it a???

OpenStudy (whpalmer4):

as for choice A: \[4^5 = 4*4*4*4*4 = (4*4)*4*4*4 = 16*4*4*4 = 16*(4*4)*4\]\[ = 16*16*4 = 256*4=1024\]

OpenStudy (anonymous):

Which number is composite? A. 17 B. 19 C. 23 D. 15

OpenStudy (whpalmer4):

what does it mean for the number to be composite?

OpenStudy (anonymous):

its 15 right???

OpenStudy (whpalmer4):

well, 15 = 3*5

OpenStudy (anonymous):

so thats it??

OpenStudy (whpalmer4):

as I said, what does it mean for the number to be composite?

OpenStudy (anonymous):

something can be multiplyed to get that number?

OpenStudy (whpalmer4):

right, it has factors other than 1 and itself. does 15 have factors other than 1 and 15? If so, it is composite. No need to ask me, you can figure it out yourself!

OpenStudy (anonymous):

Which is the prime factorization of 72? A. 2^3 • 3^3 B. 2^4 • 3^3 C. 2^3 • 3^4 D. 2^3 • 3^2

OpenStudy (anonymous):

??